Find the y-intercept and the x-intercept for the slope: y= -x^2+2x-2
Answer by ewatrrr(24785)


Hi
y= -x^2+2x-2 |Note use of ^ for squared term
y-intercept (when x = 0) Pt(0,-2)
y= -(x-1)^2 -1 |completing the square Vertex Pt(1,-1)
x-intercept (when y = 0)
-(x-1)^2 -1 =0
(x-1)^2 = -1 NO real roots, Parabola does not cross the x-axis
x = 1 ± sqrt(-1)